# Defending Against Sophisticated Poisoning Attacks with RL-based Aggregation in Federated Learning

### v1

@@ 4.1.3 Baselines
We consider four poisoning attacks in FL: explicit boosting (EB) Bhagoji et al. (2019), inner product manipulation (IPM) Xie et al. (2020), local model poisoning attack (LMP) Fang et al. (2020), and RL-based model attack (RL-attack) Li et al. (2022). IPM manipulates the attacker’s gradients to ensure the inner product with true gradients becomes negative during aggregation. LMP generates malicious model updates by solving an optimization problem in each FL epoch. EB generates malicious updates through explicit enhancement, optimizing for a malicious objective designed to induce targeted misclassification. RL-attack adaptively generates attacks on the FL system using RL.
4.1.4 Parameter Settings
We train diverse global models on distinct datasets to showcase the versatility of AdaAggRL. Specifically, for MNIST, F-MNIST, and EMNIST, a Convolutional Neural Network (CNN) serves as the global model. In the case of Cifar10, the ResNet18 architecture He et al. (2016) is utilized as the global model.
For the FL system, there are 100 clients, denoted as $K=100$, with 20 malicious clients. The FL system trains for 500 epochs, wherein each epoch, the server randomly selects 10% of the clients to participate. Each client conducts one round of training on its local dataset with a learning rate of 0.05.
For defense strategies based on RL, given the continuous action and state spaces, we select Twin Delayed DDPG (TD3) Fujimoto et al. (2018) algorithm to train the defense policies in experiments. We configure parameters such that the policy model uses ’MlpPolicy’, the learning rate is set to $1\times 10^{-5}$ for the Adam optimizer, the batch size is 64 for each gradient update, and the discount factor $\gamma$ is set to 0.99. We maintain fixed initial models and random seeds for data sampling to ensure a fair comparison.
4.2 Defense Performances
@@
We demonstrate our experiment results on four datasets. Table 2 presents the testing accuracy of different FL aggregation methods. The results indicate that AdaAggRL achieves the desired robustness. Across different datasets and models, AdaAggRL demonstrates consistent defensive efficacy against all four attack scenarios, while other defense methods may fail under certain attacks. Particularly against RL-attack, where other defense methods face significant challenges and experience a noticeable degradation in defensive performance, AdaAggRL maintains stable effectiveness.
@@
Figure LABEL:fig:cifar10 illustrates the testing accuracy variation of the global model on Cifar10 dataset, considering four attacks. The server employs different aggregation rules. AdaAggRL outperforms in all settings, even without the need for a root dataset to guide model aggregation, unlike FLtrust. While Krum and C-Median achieve commendable accuracy against IPM, LMP, and EB attacks, AdaAggRL exhibits more stable performance after 200 epochs. Particularly against RL-attack, AdaAggRL demonstrates a distinct advantage with superior accuracy and convergence speed compared to other defense methods. AdaAggRL’s performance on other datasets is shown in Appendix C.
4.3 Ablation Studies
4.3.1 Impact of Current-history Similarity
@@
To illustrate the impact of the similarity metrics $S_{cl}$ between the current client and historical data distribution on the stability of the FL process, Figure LABEL:Scl depicts the defense performance of AdaAggRL compared to the case where $S_{cl}$ is not considered under LMP attack. We observe that considering the variations in $S_{cl}$ indeed enhances the convergence speed and reduces the oscillation amplitude of testing accuracy.
4.3.2 Impact of Current (history)-global Similarity
@@
Figure LABEL:Scg illustrates the defense performance of AdaAggRL compared to the case where $S_{cg}$ and $S_{lg}$ are not considered under EB attack. We observe that the malicious client data distribution obtained through gradient reversal may stably deviate from the normal distribution, leading to an inflated $S_{cl}$. If $S_{cg}$ and $S_{lg}$ are not considered, it can significantly degrade the defense effectiveness.
4.4 Analysis
4.4.1 Impact of the Number of Attackers
Figure LABEL:attacker_lmp and Figure LABEL:attacker_eb illustrate the testing accuracy of different FL methods under LMP and EB attacks as the proportion of malicious clients increases from 0% to 95%. We observe that, under the current attacks, both AdaAggRL and FLtrust can tolerate up to 90% of malicious clients. Although, for AdaAggRL, there is a slight decrease in testing accuracy as the proportion of malicious clients increases. In contrast, the remaining FL aggregation algorithms can only tolerate malicious clients below 30% under LMP attack. This indicates that AdaAggRL can maintain a relatively stable defensive performance even when facing a high percentage of malicious clients.
4.4.2 Impact of non-i.i.d. Degree
Table LABEL:tab3:subtab2 presents testing accuracy of various FL defense algorithms against diverse attacks on MNIST-0.5. Figure LABEL:fig:non_iid provides a comparative analysis of the performance of FL defense algorithms under different distribution conditions. Under non-i.i.d. condition (q=0.5), AdaAggRL’s advantage in faster convergence becomes more pronounced, particularly against IPM, LMP, and EB. In the case of LMP, baseline algorithms exhibit a certain degree of accuracy reduction, and under EB, the oscillation magnitude of C-Median’s testing accuracy increases. Especially against RL-attack, the impact of non-i.i.d. on baseline algorithms is significant, while AdaAggRL maintains relatively stable performance. This indicates the robustness of AdaAggRL in handling non-i.i.d. client datasets.
5 Related Work
5.1 Poisoning Attacks
Based on the attacker’s objectives, poisoning attacks can be classified into targeted poisoning attacks aiming to misclassify specific input sets Bhagoji et al. (2019); Baruch et al. (2019); Bagdasaryan et al. (2020) and untargeted attacks aimed at reducing the overall accuracy of the global model Fang et al. (2020); Xie et al. (2020); Shejwalkar and Houmansadr (2021). Current untargeted attack methods typically employ heuristic-based approaches Xie et al. (2020) or optimize myopic objectives Fang et al. (2020); Shejwalkar and Houmansadr (2021); Shejwalkar et al. (2022). Bhagoji et al. (2019) generate malicious updates through explicit enhancement, optimizing for a malicious objective strategically designed to induce targeted misclassification. Xie et al. (2020) manipulate the attacker’s gradients to ensure the inner product with the true gradients becomes negative. Fang et al. (2020) generate malicious updates by solving an optimization problem. However, these attack methods typically require local updates from benign clients or accurate global model parameters to generate significant adversarial impacts and often yield suboptimal results when robust aggregation rules are employed.
To address these deficiencies, Li et al. (2022) propose a model-based RL framework for guiding untargeted poisoning attacks in FL system. They utilize model updates from the server to approximate the server’s data distribution, subsequently employing the learned distribution as a simulator for FL environment. Based on the simulator’s behavior, they utilize RL to automatically generate effective attacks, resulting in significantly reducing the global model’s accuracy. Even when the server employs robust aggregation rules, this customized method can still maintain a high level of attack effectiveness.
5.2 Defenses for Poisoning Attacks
Current defense strategies against FL poisoning attacks can be categorized into two types: one involves designing local model update evaluation mechanisms to identify malicious client-submitted model parameters Cao et al. (2020); Sattler et al. (2020); Zhang et al. (2022), and the other is based on designing novel Byzantine fault-tolerant aggregation algorithms using mathematical statistics to improve the robustness of aggregation Blanchard et al. (2017); Yin et al. (2018); Sun et al. (2019); Rajput et al. (2019); Xie et al. (2019). In evaluation mechanisms, Sattler et al. (2020) divide updates into different groups based on cosine similarity between the model parameters submitted by clients, mitigating the impact of poisoning attacks. In response to more covert poisoning attacks, some evaluation methods Cao et al. (2020); Park et al. (2021) require the server to collect a portion of clean data as a basis for validating model updates. In robust aggregation algorithms, statistical methods Blanchard et al. (2017); Yin et al. (2018); Sun et al. (2019) compare local updates and remove statistical outliers before updating the global model. For example, Blanchard et al. (2017) employ the square-distance metric to measure distances among local updates and select the local update with the minimum distance as the global model parameters. Yin et al. (2018) sort the values of parameters in all local model updates and consider the median value of each parameter as the corresponding parameter value in the global model update. Sun et al. (2019) perform gradient clipping on parameter updates before aggregation. Due to the susceptibility of statistical estimates to an outlier, existing aggregation methods cannot guarantee accuracy well and are still susceptible to local model poisoning attacks Bhagoji et al. (2019); Fang et al. (2020).
6 Conclusion
@@
In this paper, we propose AdaAggRL, an RL-based Adaptive Aggregation method, to counter sophisticated poisoning attacks. Specifically, we first utilize distribution learning to simulate clients’ data distributions. Then, we use MMD to calculate the pairwise similarity of the current local model data distribution, its historical data distribution, and the global model data distribution. Finally, we use policy learning to adaptively determine the aggregation weights based on the above similarities and the reconstruction similarity. Experiments on four real-world datasets demonstrate that AdaAggRL significantly outperforms the state-of-the-art defense model for sophisticated attacks.
Limitations & Future Work AdaAggRL assumes that the simulated data distribution obtained from poisoned updates is unstable. However, if malicious clients can launch attacks based on a stable distribution, AdaAggRL may struggle to identify malicious updates by computing the similarity of simulated distributions. Therefore, future work will focus on addressing defense strategies against such covert poisoning attacks.
@@ Appendix B Details of Experimental Settings
In each round of FL training, we reconstruct 16 images from a single model update of a sampled client to simulate data distribution. The optimization process for gradient inversion is carried out for 30 iterations. However, it’s important to note that the reconstructed images are solely used to capture the distribution characteristics and are not intended for precise image reconstruction. The limited number of iterations and reconstructed images do not suffice for accurately restoring image information. Consequently, the reconstructed images do not contain discernible image details. This approach is implemented to fulfill the privacy protection requirements of FL while minimizing computational time as much as possible. The experimental results demonstrate that, under the LMP attack, when reconstructing 16 images from the MNIST dataset, the final accuracy achieves after 10 optimization iterations of gradient inversion is 0.9554, after 20 iterations is 0.9592, after 30 iterations is 0.9636, and after 50 iterations is 0.9664. When reconstructing 32 images, the final accuracy after 30 optimization iterations is 0.9628. To strike a balance between the final performance and computational resources, we opt to train using gradient inversion with 30 iterations for reconstructing 16 images. The method for gradient inversion optimization is Adam, the learning rate is 0.05, and the reconstructed images are initialized to all-zero tensors. And the fixed parameter $\beta$ in the optimization target is set to 1e-4.
@@ Appendix C Supplementary Experimental Results
Figure LABEL:fig:fashion illustrates the test accuracy variations over FL epochs for different FL aggregation methods under four attacks (IPM, LMP, EB, and RL-attack) on F-MNIST dataset. We observe that AdaAggRL achieves higher accuracy and more stable convergence, especially when facing the more sophisticated RL-attack.
@@
Figure LABEL:fig:emnist illustrates the test accuracy variations over FL epochs for different FL aggregation methods under four attacks (IPM, LMP, EB, and RL-attack) on EMNIST dataset. We observe that AdaAggRL exhibits faster convergence under IPM, achieves higher accuracy under LMP and EB, and demonstrates particularly stable training outcomes and increased accuracy when faced with RL-attack.
@@
Figure LABEL:fig:fedavg illustrates the test accuracy variations over FL epochs for AdaAggRL and FedAvg under no attacks. We observe that in the absence of attacks, AdaAggFL and FedAvg perform similarly on MNIST and F-MNIST datasets. However, in the presence of non-i.i.d. data (MNIST-0.5), AdaAggRL achieves a slightly higher test accuracy than FedAvg after 500 rounds of training.
@@
We observe a certain degree of similarity between the historical client data distribution and the global model data distribution, as shown in Figure LABEL:fig:lg. We find that the data distribution of benign clients for historical steps is similar to the global model and remains stable, but the similarity of malicious clients is lower. This is because the data distribution of the global model represents the average state of normal data, so the distribution of benign clients should always be consistent with the global model. Instead, malicious clients do not possess this property. Though the advanced RL-attack has a high similarity with the global model due to collecting the data distribution of the server, it is still not as stable as the benign client. While considering the similarity between the current client data distribution and the global model data distribution, we also take into account the similarity between historical client data distribution and global model data distribution. By incorporating the historical performance of clients, this approach aims to capture the subtle behaviors of lurking malicious clients engaging in attacks.
@@
To illustrate the impact of the similarity metrics $S_{R}$ between the current client data distribution and historical data distribution on the stability of the federated learning process, Figure LABEL:fig:Sr depicts the defense performance of AdaAggRL compared to the case where $S_{R}$ is not considered under LMP attack. We observe that considering the variations in $S_{R}$ indeed enhances the convergence speed and reduces the oscillation amplitude of testing accuracy.

### v2

@@ Introduction
We consider four poisoning attacks in FL: explicit boosting (EB) (Bhagoji et al. 2019), inner product manipulation (IPM) (Xie, Koyejo, and Gupta 2020), local model poisoning attack (LMP) (Fang et al. 2020), and RL-based model attack (RL-attack) (Li, Sun, and Zheng 2022). IPM manipulates the attacker’s gradients to ensure the inner product with true gradients becomes negative during aggregation. LMP generates malicious model updates by solving an optimization problem in each FL epoch. EB generates malicious updates through explicit enhancement, optimizing for a malicious objective designed to induce targeted misclassification. RL-attack adaptively generates attacks on the FL system using RL.
Parameter Settings
For MNIST, F-MNIST, and EMNIST, a Convolutional Neural Network (CNN) serves as the global model. In the case of Cifar10, the ResNet18 architecture (He et al. 2016) is utilized as the global model. In FL, there are 100 clients, denoted as $K=100$, with 20 malicious clients. For defense strategies based on RL, given the continuous action and state spaces, we select Twin Delayed DDPG (TD3) (Fujimoto, Hoof, and Meger 2018) algorithm to train the defense policies in experiments [^1] Source code: https://github.com/yjEugenia/AdaAggRL.. Details on parameter determination are provided in the Appendix.
Defense Performances
EB IPM LMP RL-attack
C-Median 0.9598 0.9537 0.9329 0.5550
Clipping 0.9151 0.9654 0.1944 0.5750
FLtrust 0.9231 0.9591 0.9618 0.7300
Krum 0.9325 0.7897 0.9458 0.6875
Median 0.0981 0.9549 0.1135 0.2750
AdaAggRL 0.9659 0.9658 0.9636 0.9655
(a) CNN global model, MNIST-0.1
EB IPM LMP RL-attack
C-Median 0.9466 0.9479 0.9198 0.4250
Clipping 0.7867 0.9576 0.0986 0.4900
FLtrust 0.9488 0.9414 0.9412 0.4375
Krum 0.9274 0.7608 0.9259 0.1563
Median 0.0974 0.9448 0.1032 0.1563
AdaAggRL 0.9608 0.9617 0.9604 0.9559
(b) CNN global model, MNIST-0.5
EB IPM LMP RL-attack
C-Median 0.7935 0.8123 0.7722 0.6300
Clipping 0.7249 0.8261 0.6015 0.4600
FLtrust 0.8154 0.8344 0.8386 0.6150
Krum 0.8082 0.6610 0.7942 0.5625
Median 0.1002 0.8171 0.1001 0.0938
AdaAggRL 0.8411 0.8398 0.8400 0.8337
(c) CNN global model, F-MNIST
EB IPM LMP RL-attack
C-Median 0.8780 0.8724 0.8289 0.1857
Clipping 0.6694 0.8834 0.0008 0.1700
FLtrust 0.8618 0.8741 0.8684 0.2400
Krum 0.8309 0.5418 0.8135 0.0312
Median 0.0388 0.8716 0.4331 0.1850
AdaAggRL 0.8816 0.8805 0.8776 0.8786
(d) CNN global model, EMNIST
EB IPM LMP RL-attack
C-Median 0.7091 0.7364 0.6048 0.1150
Clipping 0.1002 0.6589 0.1388 0.0850
FLtrust 0.5786 0.6709 0.7312 0.6450
Krum 0.7238 0.7028 0.7283 0.0900
Median 0.4484 0.7309 0.1018 0.0900
AdaAggRL 0.7452 0.7497 0.7583 0.7531
(e) ResNet18 global model, Cifar10
@@
Table 2 reports the testing accuracy of various FL aggregation methods across four datasets, indicating AdaAggRL’s robustness. Across different datasets and models, AdaAggRL demonstrates consistent defensive efficacy against all four attack scenarios. Notably, AdaAggRL maintains stable effectiveness against RL-attack, where other defense methods face significant challenges and experience a noticeable degradation in defensive performance. Figure 4 illustrates the testing accuracy of the global model on Cifar10 dataset under four attacks, with different aggregation rules. AdaAggRL consistently outperforms other methods, achieving superior accuracy and convergence speed, particularly against RL-attack. Unlike FLtrust, it requires no root dataset and demonstrates more stable performance after 200 epochs, surpassing Krum and C-Median in all scenarios. AdaAggRL’s performance on other datasets is shown in Appendix.
LMP and EB attempt to optimize the objective function to make the poisoned gradients statistically inconspicuous, while RL-attack mimics the behavior of normal clients. As a result, defense methods like Median, C-Median, and Clipping, which rely solely on mean or median information through gradient clipping or selection, are prone to misjudge poisoned gradient updates. These methods may end up clipping correct updates while preserving erroneous ones.
AdaAggRL’s performance is evaluated against Feddefender and FedVal on the CIFAR-10 dataset using ResNet18 over 100 FL epochs, shown in Table 4. AdaAggRL is still significantly better than the latest baselines.
IPM LMP EB RL-attack
Feddefender 0.4711 0.4196 0.1008 0.4150
FedVal 0.6052 0.0980 0.4338 0.1392
AdaAggRL 0.6984 0.7063 0.7143 0.7106
Table 3: The accuracy of aggregation methods under attacks
Ablation Studies
Impact of Current-history Similarity
@@
To illustrate the impact of the similarity metrics $S_{cl}$ on the stability of the FL process, Figure 5(c) depicts the defense performance of AdaAggRL compared to the case where $S_{cl}$ is not considered under LMP attack. We observe that considering the variations in $S_{cl}$ indeed enhances the convergence speed and reduces the oscillation amplitude of testing accuracy.
Impact of Current (history)-global Similarity
@@
Figure 5(d) illustrates the defense performance of AdaAggRL compared to the case where $S_{cg}$ and $S_{lg}$ are not considered under EB attack. We observe that the malicious client data distribution obtained through gradient reversal may stably deviate from the normal distribution, leading to an inflated $S_{cl}$. If $S_{cg}$ and $S_{lg}$ are not considered, the defense effectiveness degrades.
Analysis
Impact of the Number of Attackers
Figures 5(a) and 5(b) show testing accuracy under LMP and EB attacks as malicious client proportion increases from 0% to 95%. Both AdaAggRL and FLtrust can tolerate up to 90% of malicious clients. AdaAggRL shows a slight accuracy decline as malicious client proportions increase, while the remaining FL aggregation algorithms can only tolerate malicious clients below 30% under LMP attack. This highlights AdaAggRL’s stability against high percentages of malicious clients.
Impact of non-i.i.d. Degree
Table 2(b) reports testing accuracy of FL defense algorithms on MNIST-0.5 under various attacks, while Figure 6 compares their performance under different distribution conditions. With non-i.i.d. data (q=0.5), baselines show reduced accuracy under LMP and increased oscillations under EB, and non-i.i.d. conditions significantly impact their performance against RL-attack. AdaAggRL demonstrates faster convergence, particularly against IPM, LMP, and EB, with minimal performance decline across attacks, demonstrating robustness to non-i.i.d. data. Under significant non-i.i.d. impact, RL can adaptively lower the weights of global model-related similarity scores. More results for q$>$0.5 can be found in the Appendix.
Conclusion
@@
In this paper, we propose AdaAggRL, an RL-based Adaptive Aggregation method, to counter sophisticated poisoning attacks. Specifically, we first utilize distribution learning to simulate clients’ data distributions. Then, we use MMD to calculate the pairwise similarity of the current local model data distribution, its historical data distribution, and the global model data distribution. Finally, we use policy learning to adaptively determine the aggregation weights based on the above similarities and the reconstruction similarity. Experiments on four real-world datasets demonstrate that AdaAggRL significantly outperforms the state-of-the-art defense model for sophisticated attacks. Future work can involve investigating novel methods to attack the adaptive defense approach. One potential scheme to construct malicious update parameters is to solve an optimization problem under the condition of controlling the stability of simulated data distribution changes.
Acknowledgments
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62406013, the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing(GJJ-23-006) and the Fundamental Research Funds for the Central Universities.
@@ Appendix B Details of Experimental Settings
The experiments are conducted on one Nvidia A100 40G GPU. We implement all models by PyTorch (Paszke et al. 2019). In FL, there are 100 clients, denoted as $K=100$, with 20 malicious clients. It trains for 500 epochs, wherein each epoch, the server randomly selects 10% of the clients to participate. Each client conducts one round of training on its local dataset with a learning rate of 0.05. For defense strategies based on RL, given the continuous action and state spaces, we select Twin Delayed DDPG (TD3) (Fujimoto, Hoof, and Meger 2018) algorithm to train the defense policies in experiments. We configure parameters such that the policy model uses ’MlpPolicy’, the learning rate is set to $1\times 10^{-5}$ for the Adam optimizer, the batch size is 64 for each gradient update, and the discount factor $\gamma$ is set to 0.99. We then present the results obtained using TD3. We maintain fixed initial models and random seeds for data sampling to ensure a fair comparison.
@@
In each round of FL training, we reconstruct 16 images from a single model update of a sampled client to simulate data distribution. The optimization process for gradient inversion is carried out for 30 iterations. However, it’s important to note that the reconstructed images are solely used to capture the distribution characteristics and are not intended for precise image reconstruction. The limited number of iterations and reconstructed images do not suffice for accurately restoring image information. Consequently, the reconstructed images do not contain discernible image details. This approach is implemented to fulfill the privacy protection requirements of FL while minimizing computational time as much as possible. The experimental results demonstrate that, under the LMP attack, when reconstructing 16 images from the MNIST dataset, the final accuracy achieves after 10 optimization iterations of gradient inversion is 0.9554, after 20 iterations is 0.9592, after 30 iterations is 0.9636, and after 50 iterations is 0.9664. When reconstructing 32 images, the final accuracy after 30 optimization iterations is 0.9628. AdaAggRL defense performance is weakly correlated with max_iters and num_images, not leading to unbounded computational overhead. Increasing max_iters to 50 only slightly improves accuracy over max_iters=10 (by about 1.15%), and num_images=16 achieves comparable accuracy to num_images=32.
To strike a balance between the final performance and computational resources, we opt to train using gradient inversion with 30 iterations for reconstructing 16 images. The method for gradient inversion optimization is Adam, the learning rate is 0.05, and the reconstructed images are initialized to all-zero tensors. And the fixed parameter $\beta$ in the optimization target is set to 1e-4.
@@ Appendix C Supplementary Experimental Results
Figure 7 illustrates the test accuracy variations over FL epochs for different FL aggregation methods under four attacks (IPM, LMP, EB, and RL-attack) on F-MNIST dataset. We observe that AdaAggRL achieves higher accuracy and more stable convergence, especially when facing the more sophisticated RL-attack.
(a) IPM
(b) LMP
(c) EB
(d) RL-attack
Figure 7: The variation in the testing accuracy of the global model over FL epochs on F-MNIST dataset considering different attacks.
@@
Figure 8 illustrates the test accuracy variations over FL epochs for different FL aggregation methods under four attacks (IPM, LMP, EB, and RL-attack) on EMNIST dataset. We observe that AdaAggRL exhibits faster convergence under IPM, achieves higher accuracy under LMP and EB, and demonstrates particularly stable training outcomes and increased accuracy when faced with RL-attack.
(a) IPM
(b) LMP
(c) EB
(d) RL-attack
Figure 8: The variation in the testing accuracy of the global model over FL epochs on EMNIST dataset considering different attacks.
Figure 8 illustrates the test accuracy variations over FL epochs for different FL aggregation methods under four attacks (IPM, LMP, EB, and RL-attack) on MNIST-0.9 (q=0.9) dataset. We observe that in the case of more extreme heterogeneity of data distribution, AdaAggRL has higher convergence speed and accuracy, and has a stronger ability to re-resume training.
(a) IPM
(b) LMP
(c) EB
(d) RL-attack
Figure 9: The variation in the testing accuracy of the global model over FL epochs on MNIST-0.9 dataset considering different attacks.
@@
Figure 10 illustrates the test accuracy variations over FL epochs for AdaAggRL and FedAvg under no attacks. We observe that in the absence of attacks, AdaAggFL and FedAvg perform similarly on MNIST and F-MNIST datasets. However, in the presence of non-i.i.d. data (MNIST-0.5), AdaAggRL achieves a slightly higher test accuracy than FedAvg after 500 rounds of training.
(a) MNIST-0.1
(b) MNIST-0.5
(c) F-MNIST
Figure 10: The comparison of test accuracy between FedAvg and AdaAggRL under attack-free conditions on different datasets.
@@
We observe a certain degree of similarity between the historical client data distribution and the global model data distribution, as shown in Figure 11. We find that the data distribution of benign clients for historical steps is similar to the global model and remains stable, but the similarity of malicious clients is lower. This is because the data distribution of the global model represents the average state of normal data, so the distribution of benign clients should always be consistent with the global model. Instead, malicious clients do not possess this property. Though the advanced RL-attack has a high similarity with the global model due to collecting the data distribution of the server, it is still not as stable as the benign client. While considering the similarity between the current client data distribution and the global model data distribution, we also take into account the similarity between historical client data distribution and global model data distribution. By incorporating the historical performance of clients, this approach aims to capture the subtle behaviors of lurking malicious clients engaging in attacks.
(a) IPM
(b) LMP
(c) EB
(d) RL-attack
Figure 11: The statistical results of the similarity between the historical client data distribution and the global model data distribution under four types of attacks vary with the training epochs on MNIST dataset.
@@
To illustrate the impact of the similarity metrics $S_{R}$ between the current client data distribution and historical data distribution on the stability of the federated learning process, Figure 12 depicts the defense performance of AdaAggRL compared to the case where $S_{R}$ is not considered under LMP attack. We observe that considering the variations in $S_{R}$ indeed enhances the convergence speed and reduces the oscillation amplitude of testing accuracy.
(a) LMP
Figure 12: The defense performance of AdaAggRL compared to the case where $S_{R}$ is not considered under LMP attack